# High-Fidelity Simultaneous Speech-To-Speech Translation

### v1

@@ 4.1 Architectural hyper-parameters
As presented in Section 3.1.4, Hibiki consists of a Temporal Transformer with a latent dimension of 2560 (7040 for the SiLU gating), 24 layers, and 128 dimensions per head, i.e. 2.2B parameters. The Depth Transformer initially follows  Défossez et al. (2024), i.e. 6 layers per codebook, a latent dimension of 1024 (2816 for the gating), and 64 dimensions per head. It models $Q=16$ audio codebooks for the output stream, and the same for the input stream (only at training). To reduce the footprint of the Depth Transformer we distill it post-training into a smaller one, with 4 layers per codebook, weight sharing for codebooks 9 to 16, low-rank embedding tables of dimension 128, and a dimension of 2048 for the gating. This reduces its size from 1.1B parameters to 449M parameters, for a total of 2.7B parameters. We also train Hibiki-M as a distilled version of Hibiki, a 1.7B variant capable of running real-time on device, with a Temporal Transformer with a latent space of dimension 2048, and 16 layers, and only 8 codebooks levels per stream.
@@ 4.2 Training protocol
Training of Hibiki-M. It goes through the same text and audio pre-training stages. During the speech translation training it is soft distilled from Hibiki, before going through the same fine-tuning step.
@@ 4.6 Results
Table 1 compares Hibiki with offline baselines that have access to the complete source audio when translating. Despite performing simultaneous translation, Hibiki outperforms all models, including the offline variant of StreamSpeech. Table 2 benchmarks Hibiki against available baselines for simultaneous translation. In the short-form setting, our model outperforms StreamSpeech and Seamless at the cost of an average 1 second of additional lagging. The long-form dataset represents a more significant challenge, as StreamSpeech does not manage to produce intelligible translations. Hibiki outperforms Seamless, again with a latency higher by an average of 1.4 seconds.
@@
No CFG 26.2 0.44
$\gamma=3.0$ (default) 25.9 0.52
$\gamma=10.0$ 22.5 0.50
@@
Ablation: Classifier-free guidance. Table 5 shows that using the very_good label provides a speaker similarity of 0.44, similar to that of Seamless (0.43). Using classifier-free guidance with $\gamma=3.0$ significantly improves it without significantly hurting translation quality, while increasing its weight too much results in degraded performance due to unintelligible speech. Supplementary material interestingly illustrates how increasing $\gamma$ to extreme values results in an exaggerated French accent (the source language in our experiments), which we can attribute to biases in the speaker model used to label our data.
@@ 4.6.1 Inference capabilities
On-device inference. Our distilled Hibiki-M reaches an 26.0 ASR-BLEU on Audio-NTREX, as shown in Table 2. This is comparable to Hibiki and better than Seamless, yet with a significantly worse speaker similarity. We attribute this weaker voice transfer ability to the lower number of quantizers modeled by Hibiki-M (8 instead of 16) which results in a twice lower audio bitrate. Figure 6 shows inference traces of Hibiki-M on an iPhone 16 Pro. Hibiki-M remains faster than real-time along a minute of inference, even with a batch size of 2 which is necessary for classifier-free guidance. Training Hibiki-M with sliding window attention would furthermore improve real-time factor along time.
@@ Appendix A Appendix
Figure 7: Example of contextual alignments. We compute the contextual alignment ($a_{j}^{\text{ctx}}$) for four different samples and plot the associated input timestamps. Some results as the two at the bottom present extreme spikes meaning output words referring to input words very far in the future. These spikes as considered as anomalies and are smoothed out as explained in Section 3.2.2.

### v2

@@ 4.1 Architectural hyper-parameters
Hibiki consists of a Temporal Transformer with a latent dimension of 2560 (7040 for the SiLU gating), 24 layers, 20 heads and local attention over 500 tokens, i.e. 2.2B parameters and a 40s context. The Depth Transformer initially follows  Défossez et al. (2024), i.e. 6 layers per codebook, a latent dimension of 1024 (2816 for the gating), and 64 dimensions per head. It models $Q=16$ audio codebooks for the output stream, and the same for the input stream (only at training). To reduce the footprint of the Depth Transformer we distill it post-training into a smaller one, with 4 layers per codebook, weight sharing for codebooks 9 to 16, low-rank embedding tables of dimension 128, and a dimension of 2048 for the gating. This reduces its size from 1.1B parameters to 449M parameters, for a total of 2.7B parameters. We also train Hibiki-M as a distilled version of Hibiki, a 1.7B variant capable of running real-time on device, with a Temporal Transformer with a latent space of dimension 2048, and 16 layers, and only 8 codebooks levels per stream.
@@ 4.2 Training protocol
Training of Hibiki-M. It goes through the same text and audio pre-training stages. During the speech translation training it is soft distilled from Hibiki, before going through the same fine-tuning step (without distillation).
@@ 4.6 Results
Table 1 compares Hibiki with offline baselines that have access to the complete source audio when translating. Despite performing simultaneous translation, Hibiki outperforms all models, including the offline variant of StreamSpeech. Table 2 benchmarks Hibiki against available baselines for simultaneous translation. In the short-form setting, our model outperforms StreamSpeech and Seamless at the cost of an average 0.7s of additional lagging. The long-form dataset represents a more significant challenge, as StreamSpeech does not manage to produce intelligible translations. Hibiki outperforms Seamless, again with a latency higher by an average of 0.8s.
@@
No CFG 26.0 0.42
$\gamma=3.0$ (default) 26.6 0.48
$\gamma=10.0$ 18.9 0.44
@@
Ablation: Classifier-free guidance. Table 5 shows that using the very_good label provides a speaker similarity of 0.42, similar to that of Seamless (0.43). Using classifier-free guidance with $\gamma=3.0$ significantly improves it without significantly hurting translation quality, while increasing its weight too much results in degraded performance due to unintelligible speech. Supplementary material interestingly illustrates how increasing $\gamma$ to extreme values results in an exaggerated French accent (the source language in our experiments), which we can attribute to biases in the speaker model used to label our data.
@@ 4.6.1 Inference capabilities
On-device inference. Our distilled Hibiki-M is competitive with Seamless on short-form and long-form translation as shown in Table 2. We attribute the lower speaker similarity on long-form audio to the lower number of quantizers modeled by Hibiki-M (8 instead of 16) which results in a twice lower audio bitrate. Figure 6 shows inference traces of Hibiki-M on an iPhone 16 Pro. Hibiki-M remains faster than real-time along a minute of inference, even with a batch size of 2 which is necessary for classifier-free guidance. Training Hibiki-M with sliding window attention would furthermore improve real-time factor along time.
@@ Appendix A Appendix
Figure 7: Example of contextual alignments. We compute the contextual alignment ($a_{j}^{\text{ctx}}$) for four different samples and plot the associated input timestamps. Some results as the two at the bottom present extreme spikes meaning output words referring to input words very far in the future. These spikes are considered as anomalies and are smoothed out as explained in Section 3.2.2.